# Particle detection performance and Geant4 simulation with low-cost CMOS technology

### v1

@@ Abstract
We analyze the performance of an Omnivision OV5647 CMOS image sensor (5 Mp) for measuring the radiation emitted from Sr90 and Cs137 sources. Our experimental arrangement includes a Raspberry Pi 3 mini-computer for data taking, processed with Python and OpenCV libraries. We specify the camera settings to be sensitive to detecting electrons and photons. We also implement a detailed Geant4 simulation of the CMOS sensor and the radioactive sources. This simulation connects the deposited energy in the pixel matrix by the electrons and photons from the radioactive sources and the ADC counts. Our measurements are expressed through the cluster size, the maximum ADC signal per cluster, and the variation of the clusters with different distances. We find a good agreement between the data and the Geant4 simulation for all these observables. Furthermore, we can reproduce the correlation between the cluster size and the maximum ADC per cluster.
keywords:
CMOS image detector, particle detection, Geant4
[…]
[inst1]organization=Sección Física, Departamento de Ciencias, Pontificia Universidad Católica del Perú,addressline=Av. Universitaria 1801, city=Lima, postcode=15088, country=Perú
[…]
[inst2]organization=UNSAAC,addressline=Av. De la Cultura, 733,, city=Cusco, postcode=8000, country=Perú
@@ 1 Introduction
Imaging sensors based on metal oxide semiconductors (CMOS), designed to take photographs, have been widely used to detect particles such as gamma rays, electrons, alphas, etc PEREZ2016171 ; 7237371 ; 8645967 ; Lipovetzky2020 . One low-cost commercial CMOS sensor that has been used for this purpose is the 5 Mp OmniVision OV5647 camera, which has shown good results in particle detection, such as discriminating between alpha and non-alpha particles by identifying and extracting ionization events 8645967 and photon imaging using fluorescence X-rays and gamma rays Lipovetzky2020 ; LIPOVETZKY2020108244 .
In this work, we define a simple experimental setup to detect electrons and photons from radioactive sources using the OV5647 sensor inside a dark box. The setup includes a Raspberry Pi 3b for data taking and precisely configuring the sensor settings with Picamera libraries, for making stable and sensitive measurements.
To reproduce the data, we implement a detailed simulation of the geometry and materials of the OV5647 sensor with Geant4 and also simulate its irradiation with radioactive sources (Sr90 and Cs137). We transform deposited energy to ADC counts and then identify clusters of pixels signaling particle events.
@@ 2 Experimental setup
To test the CMOS sensor for particle detection, we designed the experimental configuration shown in Fig.1 (devices not to scale). In the lower part, we placed a lift table, which moves the sensor along the $z$ axis and has a resolution of $100\pm 50$ $\mu$m, allowing us to control the distance between it and the radioactive source.
On the lift table, we put an aluminum dark box (68.87±$\pm 0.01$ mm high with a squared base of side 88.06±$\pm 0.01$ mm), black painted and sealed with black PVC insulation tape to ensure it is light-tight. Inside the dark box, we placed a radioactive source fixed to a support inside the box. The detector (OmniVision OV5647 CMOS sensor datasheets_rasp_camera_FWC ) is placed at the base of the dark box.
To capture the frames with the OV5647 camera we used a Raspberry Pi 3b (i.e. a small, low-cost single-board computer datasheets_raspberry3B ; raspberry ). To avoid light leakage to a minimum we connect the OV5647 camera via a flexible cable to the Raspberry Pi 3b.
@@ 2.1 OmniVision OV5647 Sensor
[…]
The CMOS sensor pixels can detect photons and electrons. In the case of photons, the photoelectric effect converts them into electron-hole pairs. Electrons ionize the medium also creating electron-hole pairs, which are later collected by the electronics. The energy used to generate an electron-hole pair in Si is on average 3.6 eV 8522044 ; BRIGIDA2004322 ; 8645967 ; SCHOLZE2000208 ; LECHNER1996206 ), for photons, as well as, approximately, for electrons. The maximum possible electrical load is called "Full Well Capacity", (FWC). Above the FWC the pixel will saturate and the accumulated charge will be equal to a maximum value 8645967 ; LANE201229 ; Bronnimann2014 . The FWC of the OV5647 sensor pixels is approximately 4300 electrons according to the datasheet datasheets_rasp_camera_FWC .
To get the source as close to the sensor as possible, the optical lens of the OmniVision OV5647 camera was removed, exposing the pixel sensor as shown in Fig. 1.
@@ 2.2 Radioactive sources
The radioactive sources that were used are Sr90 and Cs137 from Spectrum Techniques in the form of cylindrical tablets. Inside an epoxy cylinder of 0.25" in diameter with a height of 0.11" is the radioactive source. This disk, as shown in Fig. 2, is encased in a plexiglass cylinder of 1.010" diameter with a height of 0.125" spectrumtechniques .
Figure 2: Radioactive source geometry provided by the manufacturer spectrumtechniques .
The Sr90 and Cs137 sources decay almost entirely by emission of $\beta^{-}$ particles (electrons). The Sr90 source emits electrons with a maximum energy of 0.546 MeV and the resulting Y90 isotopes also decay into electrons with a maximum energy of 2.28 MeV. Thus, Sr90 is technically a pure electron source, since gamma photon emission is negligible (nds_iaea_nuclides , 10.1093/oxfordjournals.rpd.a006705 ). The characteristics provided by the manufacturer show that the Sr90 source was produced in July 2015 with an activity of 0.1 $\mu$Ci and a half-life of 28.8 years. With these data, we calculate the activity during data taking (i.e. October 2023) to be 3028 Bq.
The electron emission in Cs137 decay is about 94.6% with an energy of 0.514 MeV, producing Ba-137m, which emits a 0.662 MeV photon. Thus, Cs137 is a mixed source of electrons and photons nds_iaea_nuclides , 10.1093/oxfordjournals.rpd.a006705 . The characteristics provided by the manufacturer show that the Cs137 source was produced in June 2015 with an activity of 0.25 $\mu$Ci and a half-life of 30.1 years. With these data, we calculate the activity during data taking (i.e. October 2023) to be 0.21 $\mu$Ci.
@@ 3 Data acquisition
To record the data with our experimental setup we use a Raspberry Pi3 and Picamera libraries picamera . The latter allows us to configure the parameters of the OV5647 sensor. The regular use of the OV5647 is to take photographs, with that purpose various camera settings can change according to light conditions. Therefore, to ensure stability in the data acquisition we require to fix the following settings:
  • 1.
    Shutter speed = 0.5 seconds (i.e. the exposure time with which each frame will be taken). This time is enough to capture $\mathcal{O}(10^{3})$ radioactive decay from the selected sources.
  • 2.
@@
  • 4.
    Digital gain = 1. This is an artificial gain done by software set to its minimum value.
  • 5.
@@
For capturing images we use the 10-bit Bayer format for the ADC counts. This format is defined by three color matrices (RGB) distributed in the CMOS sensor, so that 25% of the pixels are red, 50% green, and 25% blue picamera . Since we are interested only in the intensity of each pixel and not its color, we add these three partial matrices to obtain a full matrix with the intensities in ADC values and save them in Python ’npz’ format.
To obtain the background we captured 1000 frames, equivalent to 500 seconds of actual exposure time, with the OV5647 inside the dark box without a radioactive source. The processing of each frame with the Raspberry Pi including the matrix conversion and the data saving is approximately 11.5 seconds per frame, adding to a total of about 3 hours and 11 minutes of data taking.
[…]
Additionally, with the help of the lift table (Fig.1), the detector is moved away from the radioactive source, which is fixed by the support until there is a separation of 2 mm between them, and we take 100 frames each 0.5 seconds of data every 2 mm. This procedure is repeated every 2 mm until an 18 mm separation between the source and the detector is reached. The latter has been implemented to test the inverse square distance law based only on the Sr90 source measurements.
@@ 4 Geant4 Simulation
We used a detailed Geant4 AGOSTINELLI2003250_geant4 simulation of the OmniVision OV5647 CMOS sensor to reproduce the experimental data. For this, the dimensions and materials of the CMOS sensor mentioned in section 2.1 were used to define its geometry and composition. The geometry of the radioactive sources (i.e., Sr90 and Cs137) was simulated according to the specifications provided by the manufacturer, mentioned in section 2.2. This paper’s simulation code and scripts can be found in the supplementary material scripts_data .
A fragment of the simulated CMOS sensor geometry and the simulated radioactive source geometry are shown in Fig. 3. The dark box was not simulated since it is large enough compared to the sensor, not affecting the measurement. Also, the distance between the source and the sensor is smaller than the dark box size.
@@
The number of simulated events is equivalent to the number of disintegrations for a given time, i.e., the activity of the radioactive source at the measurement time. The Sr90 source had an activity corresponding to 1514 events every 0.5 seconds (i.e., the exposure time of each frame) at the time of the measurement on October 2023. The Cs137 activity corresponds to 3811 events per 0.5 seconds.
Fig. 4 shows an example of 1 event of the Geant4 simulation of the particles emitted by a Cs137 radioactive source to the CMOS sensor, where the track left by the electron is red, the electron antineutrino track is white, and the photon is green.
@@
The simulation in Geant4 gives us the energy deposited in the pixel matrix. We convert this deposited energy into electron-hole pairs in Si, using the aforementioned factor of 3.6 eV per pair. From the number of pairs, we extrapolate the number of electrons used in the rest of the simulation chain.
After transforming the deposited energy per pixel into a number of electrons, we convert the number of electrons of each pixel into ADC counts. As we have pointed out, the maximum number of electrons before saturation is 4300 electrons, corresponding to the FWC. The minimum is 5 electrons, equivalent to one ADC count (obtained from dividing 4300 electrons over 1023 ADC counts).
Finally, we performed background measurements for three days at different times, showing a variation of 0.04% to 0.08% of active pixels, the average ADC value was between 11.5 and 11.7 with an error of 1.8, and their standard deviation was 4.2 $\pm$ 0.8. This variation is minimal, so we can consider that the background is stable over time. Fig. 5, shows the distribution of the average and standard deviation of the data taken for the background noise.

### v2

@@ Abstract
We evaluate the performance of an Omnivision OV5647 CMOS image sensor (5 Mp) for detecting radiation from Sr90 and Cs137 sources. Our experimental setup uses a Raspberry Pi 3 mini-computer for data acquisition, with image processing using Python and OpenCV libraries. We specify the necessary settings to convert a standard camera into a particle detector sensitive to electrons and photons, including a two-step background filtering procedure. In addition, we implement the first detailed Geant4 simulation that describes the layered geometry and material composition of a commercial CMOS sensor along with the radioactive sources. To enhance the simulation, we include an algorithm for charge diffusion and conversion of the energy deposited by electrons and photons into ADC counts. Our measurements are presented in terms of cluster size, the maximum ADC signal per cluster, and the number of clusters as a function of distance. We find a good agreement between the experimental data and simulation for all these observables, and we can reproduce the correlation between cluster size and maximum ADC signal per cluster. Thus, this simulation, cross-checked with data, can be used to test the feasibility of further particle detection ideas without the need to implement an experimental setup. However, the sensor has limited primary energy resolution and is thus unable to distinguish between different radioactive sources. Nevertheless, given the accurate measurement of energy deposition, the sensor, once calibrated, is suitable for dosimetric measurements of source activities.
@@ 1 Introduction
Imaging sensors based on metal oxide semiconductors (CMOS), designed to take photographs, have been widely used to detect particles such as gamma rays, electrons, alphas, etc [1, 2, 3, 4]. One low-cost commercial CMOS sensor that has been used for this purpose is the 5 Mp OmniVision OV5647 camera, which has shown good results in particle detection, such as discriminating between alpha and non-alpha particles by identifying and extracting ionization events [3] and photon imaging using fluorescence X-rays and gamma rays [4].
In this work, we define a straightforward experimental setup to detect electrons and photons from radioactive sources (Sr-90 and Cs-137) using the OV5647 sensor enclosed within a dark box. The setup includes a Raspberry Pi 3b for data acquisition. We provide a detailed description of the camera settings, configured using the Picamera libraries, to ensure stable and sensitive measurements. In addition, we outline a procedure to effectively eliminate background noise, ensuring the detection of pure signals from radioactive sources, identifying clusters of pixels corresponding to particle events.
To reproduce the experimental data, we implement, for the first time, a detailed simulation of a commercial sensor. Our simulation is divided into two main steps. First, we use Geant4 to model the geometry and materials of the sensor and radioactive sources, yielding the energy deposited in the pixel matrix. In the second step, we employ a custom algorithm to simulate charge diffusion, and to convert energy deposition into ADC counts.
@@ 2 Experimental setup
To test the CMOS sensor for particle detection, we designed the experimental configuration shown in figure1 (devices not to scale). In the lower part, we placed a lift table, which moves the sensor along the $z$ axis and has a resolution of $100\pm 50$ $\mu$m, allowing us to control the distance between it and the radioactive source.
On the lift table, we put an aluminum dark box (68.87±$\pm 0.01$ mm high with a squared base of side 88.06±$\pm 0.01$ mm), black painted and sealed with black PVC insulation tape to ensure it is light-tight. Inside the dark box, we placed a radioactive source fixed to a support inside the box. The detector (OmniVision OV5647 CMOS sensor [5]) is placed at the base of the dark box.
To capture the frames with the OV5647 camera, we used a Raspberry Pi 3b (i.e. a small, low-cost single-board computer [6, 7] ). To avoid light leakage to a minimum, we connect the OV5647 camera via a flexible cable to the Raspberry Pi 3b.
@@ 2.1 OmniVision OV5647 Sensor
The OmniVision OV5647 sensor (5 Mp) is a low-cost camera module that can be used along with a Raspberry Pi. It has a resolution of 2592 $\times$ 1944 pixels, with a pixel pitch of 1.4 $\times$ 1.4 $\mu$m2 and a 3.67 $\times$ 2.73 mm2 active area [5]. The cross-section of the OV5647 CMOS sensor, as measured in [8, 4], shows the following layers and their thickness from top to bottom: microlenses for each pixel made from PDMS (SiOC2H6) (0.735 $\mu$m), the photoresist (C10H6N2O) color filter to enable RGB imaging (0.9 $\mu$m), a thin insulator layer of SiO2 (0.225 $\mu$m) and the sensitive Si detection volume (2 $\mu$m).
CMOS sensor pixels can detect photons and electrons in different ways. The detection of visible light photons is based on the photoelectric effect, where photons create electron-hole pairs directly in silicon. The detection of electrons and high-energy photons (X-rays or gamma rays) is based on material ionization. These high-energy photons interact with the material through other processes such as photoelectric absorption or Compton scattering, releasing electrons that lose their energy as they cross the material and are able to create electron-hole pairs. In both cases, electron-hole pairs are generated, which are then picked up by the sensor’s electronics. The energy used to generate an electron-hole pair in Si is on average 3.6 eV for both photons and electrons [9, 10, 3, 11, 12].
The maximum theoretical number of electrons a CMOS sensor can store before becoming saturated is called the "Full Well Capacity" (FWC). In contrast, during normal operation, the sensor’s photodiode typically holds a lower charge to ensure linearity and reliable performance. This is referred to as the "Well Capacity" (WC), which is the number of electrons a sensor can store at a specific gain level [3, 13, 14, 15]. The OV5647 sensor has an FWC of approximately 4500 electrons per pixel. However, the WC is slightly lower, around 4300 electrons, when using an analog gain of 8 or higher, as indicated in the datasheet [5, 16].
To get the source as close to the sensor as possible, the optical lens of the OmniVision OV5647 camera was removed, exposing the pixel sensor as shown in figure 1.
@@ 2.2 Radioactive sources
The radioactive sources that were used are Sr90 and Cs137 from Spectrum Techniques in the form of cylindrical tablets. Inside an epoxy cylinder of 0.25" in diameter with a height of 0.11" is the radioactive source. This disk, as shown in figure 2, is encased in a plexiglass cylinder of 1.010" diameter with a height of 0.125" [17].
Figure 2: Radioactive source geometry according to the information provided by the manufacturer
The Sr90 and Cs137 sources decay almost entirely by the emission of $\beta^{-}$ particles (electrons). The Sr90 source emits electrons with a maximum energy of 0.546 MeV, and the resulting Y90 isotopes also decay into electrons with a maximum energy of 2.28 MeV. Thus, Sr90 is technically a pure electron source, since gamma photon emission is negligible ([18], [19]). The characteristics provided by the manufacturer show that the Sr90 source was produced in July 2015 with an activity of 0.1 $\mu$Ci and a half-life of 28.8 years. With these data, we calculate the activity during data taking (i.e. October 2023) to be 3028 Bq.
The electron emission in the Cs137 decay is approximately 94.6% with an energy of 0.514 MeV, producing Ba-137m, which emits a 0.662 MeV photon. Thus, Cs137 is a mixed source of electrons and photons [18], [19]. The characteristics provided by the manufacturer show that the Cs137 source was produced in June 2015 with an activity of 0.25 $\mu$Ci and a half-life of 30.1 years. With these data, we calculate the activity during data taking (i.e. October 2023) to be 7622 Bq.
@@ 3 Data acquisition
To record the data with our experimental setup, we use a Raspberry Pi3 and Picamera libraries [20]. The latter allows us to configure the parameters of the OV5647 sensor. The regular use of the OV5647 is to take photographs, with that purpose various camera settings can change according to light conditions. Therefore, to ensure stability in the data acquisition we require to fix the following settings:
  • Shutter speed = 0.5 seconds (i.e. the exposure time with which each frame will be taken). This time is enough to capture $\mathcal{O}(10^{3})$ radioactive decays from the selected sources.
@@
  • Digital gain = 1. This is an artificial gain performed by software set to its minimum value.
@@
To capture unprocessed images that use a color filter pattern, we used the 10-bit Bayer format of the Picamera library for the ADC counts, which does not compress the data. This format is defined by three color matrices (RGB) distributed in the CMOS sensor, so that 25% of the pixels are red, 50% green and 25% blue [20]. Since we are interested only in the intensity of each pixel and not its color, we add these three partial matrices to obtain a full matrix with the intensities in ADC values and save them in Python ’npz’ format.
To obtain the background, we captured 1000 frames, equivalent to 500 seconds of actual exposure time, with the OV5647 inside the dark box without a radioactive source. The processing of each frame with the Raspberry Pi, including the matrix conversion and the data saving is approximately 11.5 seconds per frame, adding to a total of about 3 hours and 11 minutes of data taking.
Then a radioactive source (Sr90 and Cs137, respectively) is placed on top of the OV5647 sensor at a distance of 0 mm between the source and the detector. In this configuration, the radioactive source with a circular area of 7.92 mm2 illuminates approximately 74% of the rectangular surface of the 10.02 mm2 pixel sensor. The same amount of frames and duration was taken as in the background measurement.
Additionally, with the help of the lift table (Figure1), the detector is moved away from the radioactive source, which is fixed by the support until there is a separation of 2 mm between them, and we take 100 frames each 0.5 seconds of data every 2 mm. This procedure is repeated every 2 mm until an 18 mm separation between the source and the detector is reached. The latter has been implemented to test the inverse square distance law based only on the Sr90 source measurements.
Finally, we performed background measurements for three days at different times, showing a variation of 0.04% to 0.08% of active pixels, the average ADC value was between 11.5 and 11.7 with an error of 1.8, and their standard deviation was 4.2 $\pm$ 0.8. This variation is minimal, so we can consider the background to be stable over time.
Figure 3, shows the distribution of the average and standard deviation of the data taken for the background noise.
@@ 4 Geant4 Simulation
Our simulation is divided into two parts. First, we use Geant4 [21] to model the geometry and materials of the OmniVision OV5647 CMOS sensor and radioactive sources to reproduce the experimental data. In the second part, we employ a custom algorithm to simulate charge diffusion and convert energy deposition into ADC counts.
For the first part, the dimensions and materials of the CMOS sensor mentioned in Section 2.1 were used to define its geometry and composition. The geometry of the radioactive sources (i.e. Sr90 and Cs137) was simulated according to the specifications provided by the manufacturer, mentioned in Section 2.2. This paper’s simulation code and scripts can be found in the supplementary material [22].
To simulate in Geant4, we first describe the geometry using the “DetectorConstruction” class. We define the radioactive emission source through the “PrimaryGeneratorAction” class using the “G4GeneralParticleSource” class that allows the generation of particles from the whole volume of the source. We implement the “PhysicsList” class including radioactive decays, decays, and electromagnetic standard physics. To collect the information on the simulated events, such as deposit energy, position in the particle type sensor, etc., the “RunAction” and “EventAction” classes are used, as well as the “StackingAction” and “SteppingAction” classes.
Finally, in order to change the distance between the source and the detector and other geometrical characteristics, as well as the type of radioactive source, the "DetectorMessenger" and "RunActionMessenger" classes were implemented, which allow us to create a macro to change these values.
Additionally, we made a Python script that creates the Geant4 macro for different distances and sources as well as converting the Geant4 data to ’.npz’ format to use our Python programs.
A fragment of the simulated CMOS sensor geometry and the simulated radioactive source geometry are shown in figure 4. The dark box was not simulated because it is large enough compared to the sensor, not affecting the measurement. Also, the distance between the source and the sensor is smaller than the dark box size.
@@
The number of simulated events is equivalent to the number of disintegrations for a certain time. Given the activity of the radioactive sources defined in Sec. 2.2 the corresponding number of simulated events for the Sr90 source was 1514 and for the Cs137 source 3811 events, both equivalent to an exposure of 0.5 s.
Figure 5 shows an example of 1 event of the Geant4 simulation of the particles emitted by a Cs137 radioactive source to the CMOS sensor, where the track left by the electron is red, the electron antineutrino track is white, and the photon is green.
@@
A second step in our simulation, after Geant4 has given us the energy deposited in the pixel matrix is charge diffusion in the matrix, an inherent phenomenon in the photon-to-electron conversion process in CMOS sensors, such as the OV5647.
This diffusion occurs when electrons generated by the interaction of particles from radiation (or light) are scattered before being collected by the sensor’s pixels [23, 24, 25]. This scattering causes a loss of spatial resolution, as the charge generated by a single photon can be shared by several adjacent pixels, smoothing out the edges [23, 24, 25]. We have incorporated this phenomenon, modeled with a Gaussian distribution, into the Geant4 simulation [23].
When a photon or particle hits the semiconductor material of the sensor, a cloud of electrons and holes (electron-hole pairs) is generated due to the photoelectric effect or ionization [23]. In the sensor region where there is no strong electric field (such as between the surface and the depletion region), known as the field-free region, charges can move by thermal diffusion before being collected due to interaction with the material [23, 24, 25].
We model the spatial diffusion of the deposited charge using a Gaussian function, where the number of affected pixels was limited to 9 in a 3 × 3 area, since it is a small effect. The central pixel receives most of the energy, while adjacent pixels contribute proportionally according to the Gaussian function:
$E_{\text{fraction}}(x,y)=\frac{E_{\text{dep}}}{2\pi\sigma_{\text{total}}^{2}}\exp\left(-\frac{(x-x_{0})^{2}+(y-y_{0})^{2}}{2\sigma_{\text{total}}^{2}}\right)$ (4.1)
where:
  • $E_{\text{fraction}}(x,y)$: is the fraction of energy or charge deposited in a pixel centered at $(x,y)$.
  • $E_{\text{dep}}$: is the total energy deposited in the current step.
  • $x_{0}$ and $y_{0}$: are the coordinates of the initial interaction point.
  • $\sigma_{\text{total}}$: is the total width of the Gaussian distribution and is calculated considering both the initial spread of the charge cloud and the diffusion in the field-free region [23].
The $\sigma_{\text{total}}$ is given by [24, 25]:
$\sigma_{\text{total}}=\sqrt{{\sigma_{\text{i}}}^{2}+{\sigma_{\text{ff}}}^{2}}$ (4.2)
where $\sigma_{\text{i}}$ is the initial radius of the charge cloud and $\sigma_{\text{ff}}$ is the diffusion in the field-free region.
The initial radius of the charge cloud is calculated by [23, 24, 25]:
$\sigma_{i}=k\cdot E_{\text{dep}}^{\alpha}$ (4.3)
where $k$ and $\alpha$ are empirical constants used to convert energy into a spatial range, these constants depend on factors such as the density of the material, its crystalline structure, and the properties of the medium in which the electrons move [24, 25].
The constant $k$ is related to the initial diffusion of the charge. This constant is derived from experimental studies and is specific to the material as well as the conditions under which the experiment was performed [24, 25].
The constant $\alpha$ is an empirical exponent that describes how the range of electrons changes with respect to energy. In most materials, this value lies between 1 and 2, depending on the electron energy and the characteristics of the material [24, 25].
The diffusion in the field-free region is calculated by [24, 25]:
$\sigma_{ff}=\frac{z_{ff}}{2}\sqrt{1-\left(\frac{z_{0}-z_{ff}}{z_{ff}}\right)^{2}}$ (4.4)
where $z_{ff}$ is the thickness of the field-free region, and $z_{0}$ is the z-coordinate of the interaction point.
The Gaussian diffusion parameters (i.e. $z_{ff}$, $k$, and $\alpha$) of the OV5647 CMOS sensor are not well known. To determine their most probable values, we perform a minimization search within ranges suggested by the literature and constrained by geometrical limits, comparing the simulation results with experimental data. However, since charge diffusion has only a marginal effect on the overall simulation, this minimization does not inherently guarantee agreement with the experimental data. Instead, the main source of agreement arises from the robust validation of the Geant4 simulation.
Next we define the parameters ranges. Given a total sensor thickness ($z_{max}$) of 2 $\mu$m, the field-free region ($z_{ff}$) can be between 0.2 $\mu$m and 1 $\mu$m (10-50% of $z_{max}$). Due to the absence of semiconductor doping and material properties, the empirical constants $k$ and $\alpha$ are also estimated, which were varied between 0.002 to 0.062 for $k$ and 1.73 to 1.77 for $\alpha$. Simulations covering the parameters grid are run with different values of $z_{ff}$, $k$, and $\alpha$, ensuring that the sensor dimensions are not exceeded. The grid steps for these parameters were: 0.2 for $z_{ff}$, 0.012 for $k$, and 0.005 for $\alpha$. Optimal values for these three parameters are then found by minimizing the difference between simulation and experimental measurements. In our case, the best-fit values for the OV5647 are $z_{ff}$ = 1 $\mu$m, $k$ = 0.002 $\mu$m/keV, and $\alpha$ = 1.75. The $\alpha$ value coincides with the one found for silicon in [23, 25].
After applying the charge diffusion algorithm, we convert the remaining deposited energy into electron-hole pairs in Si, using the factor of 3.6 eV (see Sec. 2.1) per pair. From the number of pairs, we extrapolate the number of electrons used in the rest of the simulation chain. After the deposited energy per pixel is transformed into a number of electrons, these electrons are converted to digital values by an Analog-to-Digital Converter (ADC) using the sensor’s WC value, which limits the maximum number of electrons a pixel can store. As we have pointed out, this number is 4300 electrons, corresponding to the WC, achieved with an analog gain of 8 which is the maximum value that the OV5647 sensor can reach to ensure stable data acquisition. Gains greater than 8 could lead to ADC saturation, as indicated in the datasheet [5, 16].
To calculate the minimum number of electrons for an ADC count, we divide the total FWC (4500 electrons) by the usable range of digital values of the ADC, which in our case is 10-bit, equivalent to a range of 1023 different values ($12^{10}-$ 1), resulting in approximately 4.4 electrons per ADC count. However, to maintain the integrity and precision of the conversion process, this value is rounded to the next integer, obtaining in this case a minimum of 5 electrons per ADC count.